OpenStudy (sarahhanish):

What is the value of this expression? (−4)4(−4)4 Enter your answer in the box.

OpenStudy (sarahhanish):

plzzz helppp

jhonyy9 (jhonyy9):

these two 4 outside parentheses are exponents ?

jhonyy9 (jhonyy9):

(-4)^4 *(-4)^4 = ? hence wrote is right ?

OpenStudy (sarahhanish):

? sorry im stupid hehe

jhonyy9 (jhonyy9):

how is this exercise in your book wrote ?

jhonyy9 (jhonyy9):

@Directrix please

OpenStudy (sarahhanish):

what you saying in wrote???

jhonyy9 (jhonyy9):

have wrote there in your book ,so how exactly ?

OpenStudy (sarahhanish):

oh wait

jhonyy9 (jhonyy9):

do you can rewriting here how is exactly in your math book please ?

OpenStudy (sarahhanish):

(-4)^4

jhonyy9 (jhonyy9):

ok. so this is the first term - yes ?

OpenStudy (sarahhanish):

yes

jhonyy9 (jhonyy9):

suppose (-4) let an (x) so than there are x^4 *x^4 = ? do you know the property of exponents for a^x *a^y = a^?

OpenStudy (sarahhanish):

uh, im not getting this

jhonyy9 (jhonyy9):

ok. a^x *a^y = a^(x+y) do you can use it in case of your exercise ?

OpenStudy (sarahhanish):

yes

jhonyy9 (jhonyy9):

so than what will get ?

jhonyy9 (jhonyy9):

suppose (-4) = a so than will be there a^4 *a^4 = a^?

Directrix (directrix):

(−4)4(−4)4 Is this: